OpenStudy (anonymous):

What is the simplest form of 2/3 - 3/8 ?

OpenStudy (anonymous):

7/24? Im not sure though.

OpenStudy (anonymous):

Find a common denominator. 2/3 - 3/8 = 16/24 - 9/24 = 7/24

OpenStudy (anonymous):

thank you very much. =)

OpenStudy (anonymous):

You're welcome :)